OpenStudy (anonymous):

Find the solution to the equation. 2x - y = -3 x + y = 0 (0, 0) (-1, 1) correct? (1, -1)

OpenStudy (anonymous):

(-1, 1)

OpenStudy (texaschic101):

2x - y = -3 -y = -2x - 3 y = 2x + 3 now sub 2x + 3 in for y in the other equation x + y = 0 x + 2x + 3 = 0 3x + 3 = 0 3x = -3 x = -1 now sub -1 in for x in either equation to find y 2x - y = -3 2(-1) - y = -3 -2 - y = -3 -y = -3 + 2 -y = -1 y = 1 check... x + y = 0 -1 + 1 = 0 0 = 0 (correct) solution is (-1,1)

OpenStudy (anonymous):

yes, thanks

OpenStudy (texaschic101):

way to go...you got the correct answer :)